OpenStudy (anonymous):

3x-4y=15 solve for y in terms of x Please solve and explain. Thank you?

jhonyy9 (jhonyy9):

add to both sides 4y 3x+4y-4y=15+4y 3x=15+4y subtract from both sides 15 3x-15=15-15+4y 3x-15=4y divide both sides by 4 (3x-15)/4 = y hope this is understandably sure easy

OpenStudy (anonymous):

Answer: Y=3/4x-15/4 Not sure how